# The thesan-zoom project: Population III star formation continues until the end of reionization

### v1

@@ Abstract
pubyear: 2025pagerange: The thesan-zoom project: Population III star formation continues until the end of reionization13
@@ 1 Introduction
Modelling Pop III star formation and feedback is notoriously challenging because minihaloes below $10^{7}{\,\rm M_{\odot}}$ require extremely high resolution (Klessen & Glover, 2023). Therefore, most of our knowledge about Pop III stars comes from small-volume “zoom-in” simulations that aim to resolve their complete formation and growth processes. Furthermore, in addition to modelling cosmic expansion, gravitational forces, hydrodynamics, and cooling, other physical ingredients are needed. For example, magnetic fields may play a role by reducing the fragmentation (Peters et al., 2014; Sharda et al., 2021; Prole et al., 2022b; Stacy et al., 2022; Saad et al., 2022; Sharda & Menon, 2024; Sadanari et al., 2024), though they can be weakened by nonideal effects (McKee et al., 2020; Sadanari et al., 2023; Mayer et al., 2025). Detailed chemical networks that also include deuterium and H-(Glover, 2005; Klessen & Glover, 2023; Nishijima et al., 2024), are necessary. One may also need to take into account the streaming velocity between dark matter and baryonic matter in the early Universe (Tseliakhovich & Hirata, 2010; Greif et al., 2011a; Schauer et al., 2019, 2021; Lake et al., 2024) and radiation transport that includes the Lyman-Werner band (Haiman et al., 1997; Skinner & Wise, 2020; Kulkarni et al., 2021; Schauer et al., 2021; Patrick et al., 2023; Sugimura et al., 2024). To determine the final IMF, these simulations must be run until the accretion stops due to internal feedback processes (Chiaki & Yoshida, 2022). These studies find that for gas-dominated cooling, a critical metallicity of $10^{-3}$ to $10^{-4}\,\text{Z}_{\odot}$ is necessary to achieve a modern universal IMF (Omukai, 2000; Bromm et al., 2001; Maio et al., 2010, 2011). In contrast, for the dust-driven transition favoured by current observations, a metallicity of $10^{-6}-10^{-4}\,\text{Z}_{\odot}$ is sufficient for producing Pop II stars (Schneider et al., 2002, 2006, 2012a; Schneider et al., 2012b; Omukai et al., 2005; Chiaki et al., 2014). Nevertheless, there are indications that the IMF only shifts toward the present IMF when metallicity exceeds $Z>10^{-2}\,\text{Z}_{\odot}$ and may also depend on the redshift due to heating from the CMB (Chon et al., 2021; Chon et al., 2022, 2024).
@@
Large-scale cosmological simulations can complement small-scale simulations, which are used to study the IMF and formation of individual Pop III stars. These larger simulations can model the full Epoch of Reionization (EoR) of the IGM ending at redshifts of $z\approx 5-6$ (Fan et al., 2006; Becker et al., 2015; Bosman et al., 2022). They also allow for a statistical analysis of the density of Pop III star formation as a function of redshift. However, these larger simulations typically have significantly worse resolution, preventing them from resolving the smallest minihaloes that are capable of forming Pop III stars. Additionally, only some simulations include an explicit Pop III subgrid model, which is typically based on the metallicity of newly formed star particles, while others apply the same feedback prescriptions used for Pop II stars. These subgrid models should be viewed as best efforts and still require calibration due to insufficient resolution. Furthermore, many studies do not use radiative transfer for the Lyman-Werner band or include molecular hydrogen chemistry, reducing their predictability. Examples of studies that analyse the formation of low metallicity stars include THESAN-HR (Borrow et al., 2023; Shen et al., 2024), simulations with dustyGadget (Di Cesare et al., 2023; Venditti et al., 2023; Venditti et al., 2024a), FLARE (Lovell et al., 2021; Vijayan et al., 2021; Wilkins et al., 2023), the Renaissance simulations (O’Shea et al., 2015; Xu et al., 2016), the GIZMO simulations from Jaacks et al. (2019) and Liu & Bromm (2020b), and the RAMSES simulations from Pallottini et al. (2014), Sarmento et al. (2018) and Sarmento & Scannapieco (2022). Almost all of these simulations predict the formation of Pop III stars up to the end of the EoR. Interestingly, the TNG-50 simulation (Pillepich et al., 2018b; Nelson et al., 2019) predicts the formation of Pop III stars even down to a redshift of $z=0$ (Pakmor et al., 2022). However, as the resolution increases, the number of star-forming haloes that consist of purely pristine gas at $z=0$ decreases, suggesting that this effect may completely disappear at sufficiently high resolutions and is an artefact of the effective equation of state used in the study.
@@
This paper is part of the thesan-zoom project (introduced in Kannan et al., 2025), which consists of zoom-in simulations of fourteen haloes spanning the halo mass range from $10^{8}\,\mathrm{M}_{\odot}$ to $10^{13}\,\mathrm{M}_{\odot}$ at the final redshift $z=3$. It includes radiative transfer with seven frequency bands, including Lyman-Werner radiation as well as the large-scale background radiation field from the parent thesan-1 box (Kannan et al., 2022; Smith et al., 2022; Garaldi et al., 2022, 2024) as boundary condition at the edge of the high-resolution region. Additionally, it includes dust and molecular hydrogen chemistry, and its dark matter resolution with particle masses down to $762\,\mathrm{M}_{\odot}$ per resolution element allows us to resolve minihaloes of $10^{6}\,\mathrm{M}_{\odot}$ by at least 1,000 particles. This corresponds to the smallest haloes capable of cooling and forming Pop III stars (Glover, 2012). While we do not implement specialized Pop III feedback or an explicit top-heavy IMF, our criterion that purely pristine gas forms Pop III stars captures first-order aspects of primordial star formation in a cosmological context, making thesan-zoom particularly well-suited for analyzing the formation of Pop III stars during the epoch of reionization.
@@ 2.1 The thesan-zoom project
Figure 1: The dark matter density (left), gas density (middle), and gas metallicity (right) centred around the target halo m10.8_8x at redshift $z=7$. We show volume averaged quantities in a slice in z-direction of depth $\rm 50\,ckpc$ and the comoving virial radius at $z=7$ (circle with solid line) and for comparison the comoving virial radius at $z=3$ (circle with dashed line). The star symbols represent Pop III star particles that formed between $z=7$ and $z=8$. All Pop III stars formed outside of the central halo but can merge later with it. The gas density around the central halo almost reaches solar metallicity, but there are still pockets of metal-free gas outside of the halo.
@@
Table 1: The main simulations we analyze in this paper: From left to right, the columns indicate the name of the simulation, the mass of the target halo at $z=3$ in the dark matter only simulation thesan-dark-1, the mass of the high-resolution dark matter and gas particles and the (minimum) softening length of (gas) star and dark matter particles. We show in appendix A results for eleven additional simulations that qualitatively agree with the three simulations discussed in the main part of this paper.
@@
For star formation and stellar feedback through stellar radiation, stellar winds and supernovae explosions, we use a significantly evolved version of the SMUGGLE model (Marinacci et al., 2019). To be eligible for star formation, a gas cell must be smaller than its thermal Jeans length and have a density greater than $n_{\mathrm{H}}=10~{}\mathrm{cm}^{-3}$. From cells meeting both criteria we generate collisionless stellar particles using a standard probabilistic method with a star formation efficiency of 100% per free-fall time. This high value prevents unresolved gas from collapsing to artificially high densities (Hu et al., 2023). Cells with mass below twice the target gas mass $m_{\rm gas}$ are fully converted into one stellar particle; otherwise, only $2m_{\rm gas}$ is removed from the gas cell to form the new particle. We assume a Chabrier Initial Mass Function (IMF, Chabrier, 2003) with minimum and maximum masses of $0.1~{}\mathrm{M}_{\odot}$ and $100~{}\mathrm{M}_{\odot}$. Age and metal-dependent luminosities are calculated using the Binary Population and Spectral Synthesis models (BPASS; Eldridge et al. 2017). The radiation is injected in the 16 nearest gas cells across seven frequency bins, without tracking the spectral shape within individual bins. Instead, averaged properties such as ionization cross-sections per bin are calculated using a $2$ Myr spectrum at quarter solar metallicity. To reduce the impact of underresolved H ii regions, we apply the corrections described in Deng et al. (2024). For the mass loss rate and energy of stellar winds from young OB stars and AGB stars, we employ analytic prescriptions from Hopkins et al. (2018), which are based on the Starburst99 (Leitherer et al., 1999) stellar evolution model. Metal enrichment rates follow Vogelsberger et al. (2013) though we note that there is a floor of $Z=10^{-8}$ in the initial conditions. Stars with masses greater than $8{\,\rm M_{\odot}}$ explode as supernovae, injecting $10^{51}~{}\mathrm{ergs}$ of energy into the ISM. To avoid the overcooling problem (Katz et al., 1996; Dalla Vecchia & Schaye, 2012) caused by not resolving the Sedov-Taylor phase, we instead inject the terminal momentum at the transition to the momentum-conserving phase (Marinacci et al., 2019). The time of each supernova event is determined stochastically based on the IMF. To better match the stellar-mass-halo-mass relations at high redshift (e.g. Moster et al., 2018; Behroozi et al., 2019), an additional feedback channel is used to inject momentum during the first $5~{}\mathrm{Myrs}$ until the first SN events occur. This ‘Early Stellar Feedback’ may represent missing physical processes and is further motivated in Kannan et al. (2025).
@@
A novel feature of the thesan-zoom project is its self-consistent modelling of the influence of external radiation sources outside the zoom-in region. This external radiation can dominate the radiation field around small haloes, leading to reduced inflows and lower star formation rates (e.g. Rees, 1986; Shapiro et al., 2004; Okamoto et al., 2008). While reionization is known to be patchy rather than uniform, most cosmological simulations typically employ a redshift-dependent, spatially uniform UV background (UVB, e.g. Faucher-Giguère et al., 2009; Haardt & Madau, 2012), combined with a purely density-based self-shielding description (e.g. Rahmati et al., 2013). As shown in Borrow et al. (2023), this neglect of the patchiness of reionization significantly affects the properties of low-mass galaxies and prevents the formation of metal-free stars at redshifts below $z=9$. In the standard simulations of the thesan-zoom project, we use the evolving UV radiation field from the parent thesan-1 simulation as a time-dependent boundary condition in the low-resolution region. Below $z=5.5$, which marks the end of the thesan-1 simulation, a smooth transition to the UVB from Faucher-Giguère et al. (2009) in the low-resolution region is implemented. Even in this case, self-shielding in the high-resolution region is done self-consistently. The thesan-1 simulation only evolved the hydrogen ionising and two helium ionising bands; therefore, our simulations do not incorporate a cosmological Lyman-Werner background, which is expected to form (Incatasciato et al., 2023), but only local sources. In a companion paper (Zier et al., 2025), we analyse in detail the effects of this improved treatment of external radiation compared to the standard uniform UVB prescription for the thesan-zoom galaxies.
@@
In this paper, we focus on the default physics runs of the 14 target haloes, which are simulated using different resolution factors, as presented in Table 1. We use the default naming scheme introduced in the introduction paper (Kannan et al., 2025), which uses the halo mass at redshift $z=3$ from the original thesan-dark-1 simulation along with the spatial zoom factor relative to the parent simulation. In the main body of this paper, we will concentrate on the simulations m12.6_4x, m10.8_8x, and m9.3_16x as representative models for different halo masses. Results from the other simulations will be shown in appendix A. We note that all results presented in the main part of the paper agree qualitatively with those presented in the Appendix.
@@ 2.2 Analysis methods and Pop III selection criteria
Figure 2: The stellar mass (within twice the stellar half mass radius) - stellar metallicity (within the stellar half mass radius) relation for the central galaxies of all target haloes at different redshifts. We always choose the highest resolution run for each halo. We compare our results with the following cosmological simulations: FiBY (Paardekooper et al., 2015), SIMBA (Davé et al., 2019), FLARES (Wilkins et al., 2023), FIRE (Ma et al., 2016), the simulation E001 from Jeong et al. (2025), MilleniumTNG (Pakmor et al., 2023; Kannan et al., 2023), and thesan-1 (Kannan et al., 2022; Garaldi et al., 2024). We also add observational results from the VANDELS survey (Cullen et al., 2019), photometry-only data from JWST (Furtak et al., 2023; Robertson et al., 2023) and spectroscopically confirmed data (Wang et al., 2023; Tacchella et al., 2023a; Curti et al., 2024; Carniani et al., 2024; Curtis-Lake et al., 2023) as well as data from dwarf galaxies within the local group (Kirby et al., 2013). Our results compare well with FiBY and the VANDELS survey. The observational results from JWST show a larger scatter than our results, though especially the spectroscopically confirmed objects agree on average with our simulation. However, they lie rather on the lower end of the metallicity. This could be explained by the typically higher redshift of the observations compared to our data points at the same stellar mass. Jeong et al. (2025) finds a higher metallicity than our results in their simulation, especially when considering their higher redshift. This could be explained by their different feedback model, especially their specific Pop III star formation model with higher metal yields. We find consistently higher metallicities than FIRE for $M_{\star}/\text{M}_{\odot}>10^{5}$. Nevertheless, the slope between FIRE and our simulations is very similar.

### v2

@@ Abstract
pubyear: 2025pagerange: The thesan-zoom project: Population III star formation continues until the end of reionization17
@@ 1 Introduction
Modelling Pop III star formation and feedback is notoriously challenging because minihaloes below $10^{7}{\,\rm M_{\odot}}$ require extremely high resolution (Klessen & Glover, 2023). Therefore, most of our knowledge about Pop III stars comes from small-volume “zoom-in” simulations that aim to resolve their complete formation and growth processes. Furthermore, in addition to modelling cosmic expansion, gravitational forces, hydrodynamics, and cooling, other physical ingredients are needed. For example, magnetic fields may play a role by reducing the fragmentation (Peters et al., 2014; Sharda et al., 2021; Prole et al., 2022b; Stacy et al., 2022; Saad et al., 2022; Sharda & Menon, 2024; Sadanari et al., 2024), though they can be weakened by nonideal effects (McKee et al., 2020; Sadanari et al., 2023; Mayer et al., 2025). Detailed chemical networks that also include deuterium and H-(Glover, 2005; Klessen & Glover, 2023; Nishijima et al., 2024), are necessary. One may also need to take into account the streaming velocity between dark matter and baryonic matter in the early Universe (Tseliakhovich & Hirata, 2010; Greif et al., 2011a; Schauer et al., 2019, 2021; Lake et al., 2024) and radiation transport that includes the Lyman-Werner band (Haiman et al., 1997; Skinner & Wise, 2020; Kulkarni et al., 2021; Schauer et al., 2021; Patrick et al., 2023; Sugimura et al., 2024). To determine the final IMF, these simulations must be run until the accretion stops due to internal feedback processes (Hosokawa et al., 2011; Sugimura et al., 2020; Chon et al., 2024). These studies find that for gas-dominated cooling, a critical metallicity of $10^{-3}$ to $10^{-4}\,\text{Z}_{\odot}$ is necessary to achieve a modern universal IMF (Omukai, 2000; Bromm et al., 2001; Maio et al., 2010, 2011). In contrast, for the dust-driven transition favoured by current observations, a metallicity of $10^{-6}-10^{-4}\,\text{Z}_{\odot}$ is sufficient for producing Pop II stars (Schneider et al., 2002, 2006, 2012a; Schneider et al., 2012b; Omukai et al., 2005; Chiaki et al., 2014). Nevertheless, there are indications that the IMF only shifts toward the present IMF when metallicity exceeds $Z>10^{-2}\,\text{Z}_{\odot}$ and may also depend on the redshift due to heating from the CMB (Chon et al., 2021; Chon et al., 2022, 2024).
@@
Large-scale cosmological simulations can complement small-scale simulations, which are used to study the IMF and formation of individual Pop III stars. These larger simulations can model the full Epoch of Reionization (EoR) of the IGM ending at redshifts of $z\approx 5-6$ (Fan et al., 2006; Becker et al., 2015; Bosman et al., 2022). They also allow for a statistical analysis of the density of Pop III star formation as a function of redshift. However, these larger simulations typically have significantly worse resolution, preventing them from resolving the smallest minihaloes that are capable of forming Pop III stars. Additionally, only some simulations include an explicit Pop III subgrid model, which is typically based on the metallicity of newly formed star particles, while others apply the same feedback prescriptions used for Pop II stars. These subgrid models should be viewed as best efforts and still require calibration due to insufficient resolution. Furthermore, many studies do not use radiative transfer for the Lyman-Werner band or include molecular hydrogen chemistry, reducing their predictability. Examples of studies that analyse the formation of low metallicity stars include THESAN-HR (Borrow et al., 2023; Shen et al., 2024), simulations with dustyGadget (Di Cesare et al., 2023; Venditti et al., 2023; Venditti et al., 2024a), FLARE (Lovell et al., 2021; Vijayan et al., 2021; Wilkins et al., 2023), the Renaissance simulations (O’Shea et al., 2015; Xu et al., 2016), the GIZMO simulations from Jaacks et al. (2019) and Liu & Bromm (2020b), and the RAMSES simulations from Pallottini et al. (2014), Sarmento et al. (2018) and Sarmento & Scannapieco (2022). Almost all of these simulations predict the formation of Pop III stars up to the end of the EoR. Interestingly, the TNG50 simulation (Pillepich et al., 2018b; Nelson et al., 2019) predicts the formation of Pop III stars even down to a redshift of $z=0$ (Pakmor et al., 2022). However, as the resolution increases, the number of star-forming haloes that consist of purely pristine gas at $z=0$ decreases, suggesting that this effect may completely disappear at sufficiently high resolutions and is an artefact of the effective equation of state used in the study.
@@
This paper is part of the thesan-zoom project (introduced in Kannan et al., 2025), which consists of zoom-in simulations of fourteen haloes spanning the halo mass range from $10^{8}\,\mathrm{M}_{\odot}$ to $10^{13}\,\mathrm{M}_{\odot}$ at the final redshift $z=3$. It includes radiative transfer with seven frequency bands, including Lyman-Werner radiation as well as the large-scale background radiation field from the parent thesan-1 box (Kannan et al., 2022; Smith et al., 2022; Garaldi et al., 2022, 2024) as boundary condition at the edge of the high-resolution region. Additionally, it includes dust and molecular hydrogen chemistry, and its dark matter resolution with particle masses down to $762\,\mathrm{M}_{\odot}$ per resolution element allows us to resolve minihaloes of $10^{6}\,\mathrm{M}_{\odot}$ by at least 1,000 particles for the smallest target halos. This corresponds to the smallest haloes capable of cooling and forming Pop III stars (Glover, 2012). For the largest target halos with a lower resolution of $4.86\times 10^{4}{\,\rm M_{\odot}}$ per dark matter particle, we still resolve larger minihalos of $10^{7}\,\mathrm{M}_{\odot}$ with 200 dark matter particles. While we do not implement specialized Pop III feedback or an explicit top-heavy IMF, our criterion that purely pristine gas forms Pop III stars captures first-order aspects of primordial star formation in a cosmological context, making thesan-zoom particularly well-suited for analyzing the formation of Pop III stars during the epoch of reionization. thesan-zoom was also used to study the star formation efficiency (Shen et al., 2025; Wang et al., 2025), the star-forming main sequence (McClymont et al., 2025a), galaxy sizes (McClymont et al., 2025b), and the impact of external reionisation on galaxy properties (Zier et al., 2025).
@@ 2.1 The thesan-zoom project
Figure 1: The dark matter density (left), gas density (middle), and gas metallicity (right) centred around the target halo m10.8_8x at redshift $z=7$. We show volume averaged quantities in a slice in z-direction of depth $\rm 50\,ckpc$ and the comoving virial radius at $z=7$ (circle with solid line) and for comparison the comoving virial radius at $z=3$ (circle with dashed line). The star symbols represent Pop III star particles that formed between $z=7$ and $z=8$. All Pop III stars formed outside of the central halo but can merge later with it. The gas around the central halo almost reaches solar metallicity, but there are still pockets of metal-free gas outside of the halo.
@@
Table 1: The main simulations we analyze in this paper: From left to right, the columns indicate the name of the simulation, the mass of the target halo at $z=3$ in the dark matter only simulation thesan-dark-1, the mass of the high-resolution dark matter and gas particles and the (minimum) softening length of (gas) star and dark matter particles. We show in appendix D results for eleven additional simulations that qualitatively agree with the three simulations discussed in the main part of this paper.
@@
For star formation and stellar feedback through stellar radiation, stellar winds and supernovae explosions, we use a significantly evolved version of the SMUGGLE model (Marinacci et al., 2019). To be eligible for star formation, a gas cell must be smaller than its thermal Jeans length and have a density greater than $n_{\mathrm{H}}=10~{}\mathrm{cm}^{-3}$. From cells meeting both criteria we generate collisionless stellar particles using a standard probabilistic method with a star formation efficiency of 100% per free-fall time. This high value prevents unresolved gas from collapsing to artificially high densities (Hu et al., 2023), but we note that the halo-scale star formation efficiency is quite robust to changes in resolution and the cell-based star formation efficiency (for a detailed discussion see Shen et al., 2025). In the ISM the criterion of being unstable according to the Jeans criterium is the more strict one, as we show in appendix A. Cells with mass below twice the target gas mass $m_{\rm gas}$ are fully converted into one stellar particle; otherwise, only $2m_{\rm gas}$ is removed from the gas cell to form the new particle. We assume a Chabrier Initial Mass Function (IMF, Chabrier, 2003) with minimum and maximum masses of $0.1~{}\mathrm{M}_{\odot}$ and $100~{}\mathrm{M}_{\odot}$. Age and metal-dependent luminosities are calculated using the Binary Population and Spectral Synthesis models (BPASS; Eldridge et al. 2017). The radiation is injected in the 16 nearest gas cells across seven frequency bins, without tracking the spectral shape within individual bins. Instead, averaged properties such as ionization cross-sections per bin are calculated using a $2$ Myr spectrum at quarter solar metallicity. To reduce the impact of underresolved H ii regions, we apply the corrections described in Deng et al. (2024). For the mass loss rate and energy of stellar winds from young OB stars and AGB stars, we employ analytic prescriptions from Hopkins et al. (2018), which are based on the Starburst99 (Leitherer et al., 1999) stellar evolution model. Metal enrichment rates follow Vogelsberger et al. (2013) though we note that there is a floor of $Z=2.7\times 10^{-9}$ in the initial conditions. Stars with masses greater than $8{\,\rm M_{\odot}}$ explode as supernovae, injecting $10^{51}~{}\mathrm{ergs}$ of energy into the ISM. To avoid the overcooling problem (Katz et al., 1996; Dalla Vecchia & Schaye, 2012) caused by not resolving the Sedov-Taylor phase, we instead inject the terminal momentum at the transition to the momentum-conserving phase (Marinacci et al., 2019). The time of each supernova event is determined stochastically based on the IMF. To better match the stellar-mass-halo-mass relations at high redshift (e.g. Moster et al., 2018; Behroozi et al., 2019), an additional feedback channel is used to inject momentum during the first $5~{}\mathrm{Myrs}$ until the first SN events occur. This ‘Early Stellar Feedback’ may represent missing physical processes and is further motivated in Kannan et al. (2025).
@@
A novel feature of the thesan-zoom project is its self-consistent modelling of the influence of external radiation sources outside the zoom-in region. This external radiation can dominate the radiation field around small haloes, leading to reduced inflows and lower star formation rates (e.g. Rees, 1986; Shapiro et al., 2004; Okamoto et al., 2008). While reionization is known to be patchy rather than uniform, most cosmological simulations typically employ a redshift-dependent, spatially uniform UV background (UVB, e.g. Faucher-Giguère et al., 2009; Haardt & Madau, 2012), combined with a purely density-based self-shielding description (e.g. Rahmati et al., 2013). As shown in Borrow et al. (2023), this neglect of the patchiness of reionization significantly affects the properties of low-mass galaxies and prevents the formation of metal-free stars at redshifts below $z=9$. In the standard simulations of the thesan-zoom project, we use the evolving UV radiation field from the parent thesan-1 simulation as a time-dependent boundary condition in the low-resolution region. Below $z=5.5$, which marks the end of the thesan-1 simulation, a smooth transition to the UVB from Faucher-Giguère et al. (2009) in the low-resolution region is implemented. Even in this case, self-shielding in the high-resolution region is done self-consistently. The thesan-1 simulation only evolved the hydrogen ionising and two helium ionising bands; therefore, our simulations do not incorporate a cosmological Lyman-Werner background, which is expected to form (Haiman et al., 2000; Incatasciato et al., 2023), but only local sources. In a companion paper (Zier et al., 2025), we analyse in detail the effects of this improved treatment of external radiation compared to the standard uniform UVB prescription for the thesan-zoom galaxies.
@@
In this paper, we focus on the default physics runs of the 14 target haloes, which are simulated using different resolution factors, as presented in Table 1. We use the default naming scheme introduced in the introduction paper (Kannan et al., 2025), which uses the halo mass at redshift $z=3$ from the original thesan-dark-1 simulation along with the spatial zoom factor relative to the parent simulation. In the main body of this paper, we will concentrate on the simulations m12.6_4x, m10.8_8x, and m9.3_16x as representative models for different halo masses. Results from the other simulations will be shown in appendix D. We note that all results presented in the main part of the paper agree qualitatively with those presented in the Appendix.
@@ 2.2 Analysis methods and Pop III selection criteria
Figure 2: The stellar mass (within twice the stellar half mass radius) - stellar metallicity (mass-weighted mean within the stellar half mass radius) relation for the central galaxies of all target haloes at different redshifts. We always choose the highest resolution run for each halo and additionally highlight the three simulations we focus on in the main part of the paper by increasing their symbol. We compare our results with the following cosmological simulations: FiBY (Paardekooper et al., 2015), SIMBA (Davé et al., 2019), FLARES (Wilkins et al., 2023), FIRE (Ma et al., 2016), the simulation E001 from Jeong et al. (2025), MilleniumTNG (Pakmor et al., 2023; Kannan et al., 2023), and thesan-1 (Kannan et al., 2022; Garaldi et al., 2024). We also add observational results from the VANDELS survey (Cullen et al., 2019), photometry-only data from JWST (Furtak et al., 2023; Robertson et al., 2023) and spectroscopically confirmed data (Wang et al., 2023; Tacchella et al., 2023a; Curti et al., 2024; Carniani et al., 2024; Curtis-Lake et al., 2023) as well as data from dwarf galaxies within the local group (Kirby et al., 2013). Our results compare well with FiBY and the VANDELS survey. The observational results from JWST show a larger scatter than our results, though especially the spectroscopically confirmed objects agree on average with our simulation. However, they lie rather on the lower end of the metallicity. This could be explained by the typically higher redshift of the observations compared to our data points at the same stellar mass. Jeong et al. (2025) finds a higher metallicity than our results in their simulation, especially when considering their higher redshift. This could be explained by their different feedback model, especially their specific Pop III star formation model with higher metal yields. We find consistently higher metallicities than FIRE for $M_{\star}/\text{M}_{\odot}>10^{5}$. Nevertheless, the slope between FIRE and our simulations is very similar.